The number of patients Dr. Wong can see is directly proportional to the time it takes her to see them. If she can see 9 patients in 3 hours, how long would it take her to see 45 patients?

Answer by rfer(16322)

45/3=15 hrs
------------
3/9=x/45
9x=135
x=15 hrs